OpenStudy (anonymous):

CAN SOMEONE PLEASE EXPLAIN THIS?!?! Find the derivative of f(x) =-9/x at x = 6.

jimthompson5910 (jim_thompson5910):

Are you familiar with the power rule?

OpenStudy (anonymous):

No I don't understand any of this I just want to finish this class so I can graduate.... @jim_thompson5910

jimthompson5910 (jim_thompson5910):

is this for calculus class or another kind of class?

OpenStudy (anonymous):

it's pre-calc @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok you can type in "derivative of f(x) =-9/x" into wolfram alpha to get https://www.wolframalpha.com/input/?i=derivative+of+f%28x%29+%3D-9%2Fx what does it say for f ' (x) ?

OpenStudy (anonymous):

are you talking about the 9/x^2 or the 9?

jimthompson5910 (jim_thompson5910):

yes 9/x^2

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

\[\Large f \ ' (x) = \frac{9}{x^2}\] \[\Large f \ ' (6) = \ ??\]

OpenStudy (anonymous):

sorry i lost connection i'm back now! was it 9/36??

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes, now reduce

OpenStudy (anonymous):

1/4?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

yay! i have one more, can you help me?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

its a short response question; The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative.

jimthompson5910 (jim_thompson5910):

you can use wolfram alpha to compute the derivative function

jimthompson5910 (jim_thompson5910):

it's unfair how they are asking you derivative questions in a pre-calc class

OpenStudy (anonymous):

i'm in the last chapter so idk

jimthompson5910 (jim_thompson5910):

what is s ' (t) ?

OpenStudy (anonymous):

to be completely honest i don't know...

jimthompson5910 (jim_thompson5910):

what do you get when you type in "derivative of s(t) = -8 - 9t" into wolfram

OpenStudy (anonymous):

-9 :)

jimthompson5910 (jim_thompson5910):

yep, in this case the derivative is the same no matter what t is

jimthompson5910 (jim_thompson5910):

basically the rate of change is the same throughout

OpenStudy (anonymous):

alright so to get credit for the short essay what else do i need to put??

jimthompson5910 (jim_thompson5910):

you can throw in how it's a linear equation which has the same slope throughout

jimthompson5910 (jim_thompson5910):

slope of linear equation = rate of change

OpenStudy (anonymous):

it's a linear equation so the rate or change remains the same, which is -9??

OpenStudy (anonymous):

how does that sound?

jimthompson5910 (jim_thompson5910):

that sounds good

jimthompson5910 (jim_thompson5910):

oh it's "rate OF change" not rate or change

OpenStudy (anonymous):

okay thanks!! :D You've been a great help!!!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

oh typo!! X)